1.If sin = 12, and is in QI, find sec data .



2.Given cos =35 , is in QIV, find csc data.


3.Find the exact values of the circular functions whose =-3π/4.


1
Expert's answer
2022-05-18T09:41:57-0400

1. As it is in Q1, then all functions are positive.

"sin \alpha =1/2"

"cos^2\alpha=1-sin^2\alpha=1-1/4=3/4"

"cos \alpha =\sqrt {3/4}=\sqrt{3}/2"

"sec\alpha=1/cos \alpha=1/(\sqrt{3}/2)=2/\sqrt3"

2. As it is in QIV, then sin<0 and csc<0

functions are positive.

"cos^2\alpha=3/5"

"sin^2\alpha=1-cos^2\alpha=1-9/25=16/25"

"sin \alpha=\sqrt{16/25}=-4/5"

"csc=1/sin\alpha=1/(-4/5)=-5/4"

3. "3\pi/4=135°=3x45°"

Of course 30/60/90 and 45/45/90 are the only two triangles students are expected to know "exactly." We have to know the multiples of these triangles in the other quadrant. Here we have 45/45/90 in the second quadrant.


We start from cos (45°)=sin (45°)="1/\sqrt2"

The angle

135∘

is supplementary to

45∘

, so has the opposite cosine and the same sine.

"cos({3\pi}/4)= - cos(\pi - {3\pi}/4)=- cos(\pi/4)= - 1/\sqrt{2}"

"sin({3\pi}/4)=sin( \pi - {3\pi}/4) = sin(\pi/4) = 1/\sqrt{2}"

"tan({3\pi}/4) = {sin({3\pi}/4)}/{cos({3\pi}/4)} = {(1/\sqrt{2})}/{(-1/\sqrt{2})}=-1"


"sec({3\pi}/4)=1/cos({3\pi}/4) = - \sqrt{2}"

"csc({3\pi}/4)=1/sin({3\pi}/4) = \sqrt{2}"

"cot({3\pi}/4)=1/tan({3\pi}/4) = -1"